Circular No. 6898

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bt L. Germany reports that the Mount Stromlo Abell Cluster Supernova

Search Team (cf. IAUC 6639, 6893) has discovered an apparent supernova (V = 19.7), presumably associated with the Abell cluster 1736, on V and R CCD images taken on Mar. 10 with the Mount Stromlo 1.27-m telescope (+ Macho Camera). No background host galaxy has been identified for the object, which is located at α = 13^{h}25^{m}41^{s}.93, δ = −26^{o}46′55′′.7 (equinox 2000.0). SN 1998bt is present on images taken on Feb. 23 and Apr. 6, 18, and 30, with limiting V magnitudes of 21.0, 21.0, 21.5, and 21.5, respectively. The object increased in brightness by > 1.5 mag between Feb. 23 and Apr. 6, and has faded by about half a magnitude since then. The star was not present on an image taken on Jan. 27 (limiting mag V = 21.5).

SUPERNOVA 1998Z IN ANONYMOUS GALAXY Germany also writes: “I obtained a spectrum (range 380–750 nm)

of SN 1998Z (cf. IAUC 6850) with the Mount Stromlo 2.3-m telescope (+ DBS) on Apr. 1. The spectrum confirms this object to be a type-Ia supernova within a week of maximum light.”

GRB 980425 P. M. Vreeswijk, University of Amsterdam, writes that he provided the

wrong position for the optical position of the suspected supernova in ESO 184-G82 on IAUC 6895, lines 7–8; the correct position should read α = 19^{h}35^{m}03^{s}.17, δ = −52^{o}50′46′′.1 (equinox 2000.0).

SAX J1808.4–3658 = XTE J1808–369 A. V. Filippenko adds that the spectra reported on IAUC 6897 were

obtained on May 2 UT.

SUPERNOVA 1998aq IN NGC 3982 CCD V photometry: Apr. 20.904 UT, 12.67 (D. Hanžl, Brno, Czech

Republic); 25.911, 12.25 (Hanžl); 26.957, 12.22 (Hanžl); 27.15, 12.14 ± 0.14 (D. B. Caton, Dark Sky Observatory, Appalachian State University); 28.027, 12.18 (Hanžl); 30.886, 12.20 (Hanžl); May 2.956, 12.23 (Hanžl); 5.947, 12.36 (Hanžl). On Apr. 20.904, Hanžl also reported B − V = +0.02.

1998 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green